OpenStudy (anonymous):

Simplify the expression using the properties of exponents: (6a^2b^3)^3 _____________ (b^2)^3

OpenStudy (anonymous):

6a^(2*3) * 2b^(3*3) ------------------- = b^(2*3) 6a^6 * 2b^9 ------------ b^6

OpenStudy (anonymous):

6a^6 _________ 2b^3 Is this correct??

OpenStudy (anonymous):

No, it is 6a^6 * 2b^9 ------------ and that equals 12a^6 * 2b^3 b^6

OpenStudy (anonymous):

\[\frac{(6a{}^{\wedge}2b{}^{\wedge}3){}^{\wedge}3}{(b{}^{\wedge}2){}^{\wedge}3}\text{= }216 a^6 b^3\]If the LHS and the RHS are each evaluated with a = 3 and b = 5 then 19683000 = 19683000 is the result.